OpenStudy (anonymous):

please help and check my answers. thank you. log b 5 = 1.1610 log b 2 = .5000 Find each of the following. (a) log b10 log b5+ log b5= 1.661 (b)log b125 2log b5+log b5 (1.1610)^2+1.1610=2.508921 (c)log b2/5 log b2-log b5 .5000-1.1610=-.661 (d) log b3√25 completely stuck on d

OpenStudy (anonymous):

Your syntax is all smashed together, and difficult to understand. Did you mean this initially? \(\huge \log_b5=1.1610\) \(\huge \log_b2=0.5\) In which case you'd want to find what b is... \(\huge b^{\log_b 2} = b^{0.5}\) \(\huge 2 = b^{0.5}\) \(\huge 2 = \sqrt{b}\) \(\huge 4 = b\) Check: \(\huge \log_45=1.160964047...\)

OpenStudy (anonymous):

For your last one, isn't \(3\sqrt{25}\) just = 3*5 = 15? ;-) Unless that's not what you meant to write.